40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Oct 2 02:33:32 2011, kiwamu, Update, LSC, something funny with AS55 outofloop.png
    Reply  Thu Oct 6 05:18:20 2011, kiwamu, Update, LSC, Noise in AS55 was from clipping : fixed  outofloop.png
Message ID: 5624     Entry time: Thu Oct 6 05:18:20 2011     In reply to: 5595
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: Noise in AS55 was from clipping : fixed  

It turned out the noise in AS55 was due to a clipping.  After fixing the clipping the noise successfully went down.

I was going to briefly check the clipping and go ahead locking DRMI, but for some reason I couldn't stop myself from working on this issue.

Here is a plot of the noise spectra taken before and after fixing the clipping.

The configuration of this measurement is exactly the same as that I did before (#5595)

outofloop.png

(what I did)

 + Locked power-recycled ITMY so that the AS beam is bright enough to work with.

 + Shook BS at 1 Hz in the YAW direction

 + Looked around the AP table with an IR viewer and searched for a clipping moving at 1Hz.

 + Found the first lens in the AS beam path has clipped the beam at the upper side. A tiny portion of the beam was clipped.

 + Corrected the beam height to 4 inch by steering the very first mirror.

 + Raised the height of the lens because it was about 3.5 inch or so.

 + Found the lens had a scratch (~1 mm size ) at 1 cm blow the center on the surface.

   => I tried finding a spare 2 inch lens with a long focul length, but I couldn't find it,

        So I left the lens as it is, but we should buy some 2 inch lenses just in case like this.

 + Replaced the 1 inch beam splitter by 2-inch 99% BS so that most of the light goes into the RFPD and a little bit goes into the camera.

Quote from #5595

The AS55 signal contains more noise than the REFL signals.

ELOG V3.1.3-